OpenStudy (anonymous):

You roll 1 dice. Find p(the number rolled is not odd) Is it 1/2.

OpenStudy (anonymous):

Correct. Probability of it being odd is 3/6 (#odds / total) = 1/2. Probability of it not being odd is 1 - 1/2 = 1/2.